64kstxrz:

Adenan makes a scale drawing of a rectangular flower box. The length is 5 in., and the width is 3 in. Adenan changes the scale of the drawing from 1 in. : 4 ft to 1 in. : 6 ft. Which statement about the dimensions of the flower box is true? The length of the flower box under the new scale is 15 ft. The length of the flower box under the old scale is 30 ft. The width of the flower box under the old scale is 12 ft. The width of the flower box under the new scale is 24 ft.

XlilxmoonX:

is this a test or quiz

jhonyy9:

so in the first scale suing the rule of 3 simple you can calcule that when 1 in mean 4 ft 5 in mean x ft so cross multiplie and calcule the x value x = 5*4/1 = 20 ft using this way to the new scale drawing of 1 to 6 when 1 in mean 6 ft so 5 in mean x ft x = 5*6/1 = ? ft hope this is an understandably way so in this way you can calcule the width too hope helped